OpenStudy (anonymous):

PLEASE HELP!! The number of milligrams of a drug that remains in a patient's system after t hours is given by the function A(t) =ie^rt. Jordan was given 350 milligrams of medicine which leaves his bloodstream at a rate of 20%. How much of the medicine remains in his system after 4 hours?

hartnn (hartnn):

A(t) = lert ? what does that mean ?

OpenStudy (anonymous):

Ie^rt

hartnn (hartnn):

r i think is rate, what is le ?

hartnn (hartnn):

ok, wait, e is exponential, so, we need 'i' first. when A(t) = 350 350 = i e^ (20/100 * t)

hartnn (hartnn):

is it like \(\large ie^{rt}\) or \(\large ie^rt\)

hartnn (hartnn):

anyways with given information, i don't think its possible to find what is required..

OpenStudy (anonymous):

Yes it is like how you wrote it! @hartnn

hartnn (hartnn):

means ? and thats because we won't be able to find the value of i .

hartnn (hartnn):

yeah, i would say, the given information is insufficient. but if you can, post the screenshot of the question ?

OpenStudy (anonymous):

It looks just like I put it but with the answers The number of milligrams of a drug that remains in a patient's system after t hours is given by the function A(t) =ie^rt. Jordan was given 350 milligrams of medicine which leaves his bloodstream at a rate of 20%. How much of the medicine remains in his system after 4 hours? 12.91 mg 35.77 mg 108.46 mg 157.27 mg

hartnn (hartnn):

ok, then i assume, 350 mg was given 'initially' , now it'll be solvable , so, at t=0, A(t) =350, implies i=350 at t= 4 hrs, A(t) = 350 * e^ (20/100 * 4) = 350 * e^(4/5) =.... ?

OpenStudy (anonymous):

157.27 mg?

hartnn (hartnn):

i am getting 778mg which is not in the choices, how you got 157 ?

OpenStudy (anonymous):

I guessed.

OpenStudy (anonymous):

I don't understand this question!!

hartnn (hartnn):

that 'i' is bugging me. and what i get is not in the choices.....sorry.

OpenStudy (anonymous):

It's fine..Thanks for the help anyways :D Which one would you guess?! Ha

hartnn (hartnn):

guesses could be wrong, i closed my eyes and picked up C.

hartnn (hartnn):

ok, yes.

hartnn (hartnn):

oh, did you miss the -*minus* sign ??? \(\huge A(t)=ie^{-rt}\) ?

hartnn (hartnn):

yes, you did ! 350 e^(-20/100 * 4) = 157.26 -_-

OpenStudy (anonymous):

Oh no it's just ie^rt so 157.26 must be wrong..

hartnn (hartnn):

if its just ie^rt in you question, then the QUESTION must be incorrect.

OpenStudy (anonymous):

I didn't miss any signs. I'm going to email my teacher..

hartnn (hartnn):

okay.